OpenStudy (anonymous):

Solve -5y-9=-(y-1) will medal and fan :3

pooja195 (pooja195):

Start by distributing the negative sign \[\huge~\rm~\bf~−5y−9=\color{red}{−}\color{blue}{(y−1) }\]

pooja195 (pooja195):

So distribute the - into the blue (-1)(y) = ? (-1)(-1)=?

OpenStudy (anonymous):

0?

pooja195 (pooja195):

We are multiplying So distribute the - into the blue (-1)(y) = -y (-1)(-1)=1 (negative times a negative makes a postive) Now we have \[\huge~\rm~\bf~−5y−9=−y+1 \] Any ideas on what we should do next?

OpenStudy (anonymous):

Oh u.u should we combine like terms next?

pooja195 (pooja195):

correct, add y to both sides \[\huge~\rm~\bf~\color{red}{−5y−9+y=−y+1+y}\]

pooja195 (pooja195):

What is the equation we are left with?

OpenStudy (anonymous):

-5y-9y

pooja195 (pooja195):

we dont combine y with 9 we combine it with -5y because they are like terms -5+1=?

OpenStudy (anonymous):

-4

pooja195 (pooja195):

Good so now we have \[\huge~\rm~\bf~−4y−9=1 \] now add 9 to both sides \[\huge~\rm~\bf~−4y−9+9=1 +9 \] what equation are we left with ?

OpenStudy (anonymous):

-4y=10 ? :3 thank you for helping me by the way. It means a lot

pooja195 (pooja195):

perfect and yw ^_^ Now divide both sides by 4 to isolate the y y=?

OpenStudy (anonymous):

I got -2.5 n.n

pooja195 (pooja195):

perfect but leave it as a fraction \[\huge~\rm~\bf~\frac{ 10 }{ -4 }\rightarrow \frac{ ? }{ ? }\] simplfy the fraction

pooja195 (pooja195):

What is the GCF of 10 and 4?

pooja195 (pooja195):

No it does not.

OpenStudy (anonymous):

5??

OpenStudy (anonymous):

Ohhh wait is it 10? ; n ;

pooja195 (pooja195):

What can we divide 10 and 4 by to get the fraction to it's simpliest form

OpenStudy (anonymous):

We can divide it by half

OpenStudy (dangalique):

yes it does. do the math!!! 0_0 (2) is your answer. and the computer just told me i was right!!! 0_0

pooja195 (pooja195):

Correct so 2 \[\huge~\rm~\bf~\frac{ 10/2= }{ -4/2=}\rightarrow \frac{ ? }{ ? }\]

pooja195 (pooja195):

2 is incorrect. @dangalique

OpenStudy (anonymous):

5/-2

OpenStudy (dangalique):

thankyou!!! : )

pooja195 (pooja195):

Good however it would be this: \[\huge~\rm~\bf~\frac{ 10 }{ -4 }\rightarrow \frac{ -5 }{ 2 }\]

OpenStudy (anonymous):

Thank you pooja :3

pooja195 (pooja195):

yw ^_^

OpenStudy (dangalique):

that's an improper fraction. so that's not how you write it!!! 0_0

pooja195 (pooja195):

The equation does not specify what form they want it in. They simply want to isolate y please go review the material before trying to prove others wrong :P

OpenStudy (dangalique):

well in high school you dont leave your answers in improper fractions!!! 0_0